# From Long Videos to Engaging Clips: A Human-Inspired Video Editing Framework with Multimodal Narrative Understanding

### v1

@@ Abstract
The rapid growth of online video content, especially on short video platforms, has created a growing demand for efficient video editing techniques that can condense long-form videos into concise and engaging clips. Existing automatic editing methods predominantly rely on textual cues from ASR transcripts and end-to-end segment selection, often neglecting the rich visual context and leading to incoherent outputs. In this paper, we propose a human-inspired automatic video editing framework (HIVE) that leverages multimodal narrative understanding to address these limitations. Our approach incorporates character extraction, dialogue analysis, and narrative summarization through multimodal large language models, enabling a holistic understanding of the video content. To further enhance coherence, we apply scene-level segmentation and decompose the editing process into three subtasks: highlight detection, opening/ending selection, and pruning of irrelevant content. To facilitate research in this area, we introduce DramaAD, a novel benchmark dataset comprising over 800 short drama episodes and 500 professionally edited advertisement clips. Experimental results demonstrate that our framework consistently outperforms existing baselines across both general and advertisement-oriented editing tasks, significantly narrowing the quality gap between automatic and human-edited videos.
@@ 1 Introduction
Figure 1: Distinctions between common automatic video editing methods and human editors. When editing videos, human editors primarily focus on the narrative and audience experience, while prior methods typically do not consider these aspects. Inspired by human editors, our approach similarly prioritizes the video’s narrative and audience experience.
@@
To address these challenges, we propose HIVE, an automatic video editing framework based on multimodal narrative understanding. The HIVE considers both the visual content and the underlying storyline, enabling the generation of diverse and engaging edited videos. Specifically, to achieve accurate video comprehension, we incorporate not only character extraction and dialogue analysis but also leverage MLLMs to generate narrative summaries of video content, thereby compensating for missing visual information. To enhance the coherence and viewing experience of the edited clips, we adopt an improved video scene segmentation Rao et al. (2020), instead of conventional shot detection Fu et al. (2017), to divide the video into semantically complete scenes. This ensures that clip boundaries do not interrupt ongoing dialogue or character actions.
@@
To support research in automatic video editing, we present DramaAD, a novel benchmark dataset constructed from real-world short drama videos. These short dramas, fast-paced serialized episodes typically lasting 1 to 3 minutes, have recently surged in popularity on short video platforms. In practical applications, companies are increasingly interested in repurposing such content into advertisement-style clips through automatic editing, aiming to reduce manual effort and production costs. DramaAD addresses this emerging need by offering both academic and industrial value: it comprises over 800 raw drama videos and more than 500 professionally edited advertisement clips as references. To facilitate objective evaluation, we also propose a suite of quantitative metrics covering multiple aspects of editing quality. Experimental results demonstrate that our proposed framework significantly outperforms existing baselines in both ad-specific and general video editing tasks, while closing the performance gap with human editors.
@@ 2.1 Automatic Video Editing
In recent years, automatic video editing has gained increasing attention Soe (2021); Ronfard (2021); Koorathota et al. (2021); Wang et al. (2019). Early methods primarily relied on feature extraction for video editing. Pardo et al. (2021) trained a ranking model using contrastive learning techniques, which can rank and identify plausible cuts among all possible transitions between a given pair of shots. Podlesnyy (2020) designed a data-driven editing model that treats the video editing task as a sequence decision problem, allowing the model to determine whether each shot should be retained. Hu et al. (2023) utilized a pretrained Vision-Language Model to extract editing-related representations and train an editor model by reinforcement learning. These methods lack an explicit understanding of the video content and suffer from a lack of interpretability. Recently, LLM-based editing methods have made progress. Some works Barua et al. (2025); Wang et al. (2024a) utilize LLMs to assist human editors, while other fully automated approaches, such as FunClip, leverage ASR-transcribed speech to achieve video editing.
@@ 2.2 Video Understanding
Recently, significant progress has been made in LLM-based video understanding. Methods such as VideoChat2 (Li et al., 2024) and LLaVA-Video (Lin et al., 2023) integrate video foundation models with LLMs. VideoChat2 uses a learnable Q-former inspired by BLIP2 (Li et al., 2023) for aligning visual and textual features, while LLaVA-Video employs a simple linear projection trained jointly on image-video data. Other methods, such as ChatVideo (Wang et al., 2023a) and VideoAgent (Wang et al., 2024b), incorporate LLMs within agent frameworks. ChatVideo represents videos by trajectories and their attributes to support LLM-based reasoning. VideoAgent relies on iterative analysis, dynamically selecting keyframes and using multimodal reasoning to understand videos. Furthermore, general-purpose multimodal models, such as GPT-4o (Hurst et al., 2024) and Gemini (Team et al., 2023), also demonstrate powerful visual understanding abilities applicable to video tasks. In addition, some studies focus on understanding films through audio description (Han et al., 2023), narration (Zhang et al., 2024), and storytelling (He et al., 2024), requiring finer video understanding and stronger narrative flow.
In the short-video domain, datasets such as SkyScript-100M (Tang et al., 2024) provide structured textual formats capturing scenes, emotions, and cinematography instructions, thus supporting AI-based video generation tasks. Similarly, Vript (Yang et al., 2025) creates detailed mappings between general videos and fine-grained screenplay text, enabling more precise video description and understanding.
In contrast to these works, our research focuses specifically on video semantic analysis and leverages multi-dimensional information to achieve deeper video understanding. We integrate multi-dimensional information including character attributes, interpersonal relationships, and scene context to achieve deeper video understanding and content analysis.
Figure 2: An overview of the proposed framework. The video understanding module (a) extracts video information across multiple dimensions, such as dialogues and characters, etc., and segments the video into semantically complete scenes. Consequently, we obtain a sequence of scenes along with their corresponding narrative descriptions (b). This content is then fed into the editing module (c), which identifies highlights, selects appropriate editing boundaries, and removes unnecessary scenes as human editors do (we take highlight clip 1 for example).
@@ 3.1 Multimodal Narrative Understanding
Our framework combines audiovisual models with MLLMs to convert video content into detailed textual representations, simulating human cognition. It features character extraction, dialogue analysis, character-dialogue matching, video scene segmentation and comprehensive caption. A memory module is introduced to handle long-form videos, maintaining intermediate representations and capturing inter-segment relationships.
Character Extraction. The goal of character extraction is to identify individuals in videos and build a database of their key information. While annotated data or metadata may sometimes be available, it is often missing and must be generated automatically. To address this, we employ face detection Deng et al. (2019); Zhou et al. (2023) and clustering: detecting faces in video frames, extracting facial embeddings, and clustering them into distinct identities. Additionally, we enrich character profiles by leveraging character introductions from drama narratives. We finetune an MLLM that extracts character-related textual descriptions, such as names and identities, integrating these into the character database to facilitate deeper narrative comprehension.
Dialogue Analysis. Films and short dramas often feature dense dialogues, with key plot developments driven by character conversations. To capture this, we first apply ASR technology Radford et al. (2023); Zhang et al. (2023). However, its transcripts are prone to errors from acoustic interference, background music, and character name ambiguities. To address these issues, we use optical character recognition (OCR) Bautista and Atienza (2022); Fang et al. (2021); Wang et al. (2021) to extract subtitles directly from video frames, providing a reliable reference. Leveraging LLMs, we align and correct ASR outputs based on OCR subtitles, significantly improving transcription accuracy.
@@
4:$H=(h_{1},h_{2},...,h_{k})\leftarrow L(R_{h},V)$triangleright Extract highlight clips $H$ using LLM
5:$H^{\prime}\leftarrow\text{sort}(H)$triangleright Sort highlight clips $H$ by scene score in descending order
@@
8:     $O=\{o_{1},...,o_{m}\}\leftarrow L(R_{o},h^{\prime},V)$triangleright Select all suitable opening scenes for $h^{\prime}$, $O\subseteq V$
9:     $E=\{e_{1},...,e_{n}\}\leftarrow L(R_{e},h^{\prime},V)$triangleright Select all suitable ending scenes for $h^{\prime}$, $E\subseteq V$
10:     $B\leftarrow B\cup(O\times E)$triangleright Enumerate all openings-endings pairs by cartesian product
@@
14:     $P=\{p_{1},...,p_{l}\}\leftarrow L(R_{p},V[o:e])$triangleright Get scenes to be pruned within $V[o,e]$, $P\subseteq V$
15:     $V^{\prime}\leftarrow V^{\prime}\cup\text{Cut \& Splice}(V[o:e]\setminus P)$triangleright Cut and splice remaining scenes
@@
Character-Dialogue Matching. To enable fine-grained narrative analysis, we introduce a specialized module that links identified characters to their dialogues and infers their relationships. After applying speaker diarization Horiguchi et al. (2020), dialogue segments are initially unlabeled. Simple heuristics, such as matching visible characters to active speakers, are used but often fail in complex cases like off-screen dialogues. To enhance accuracy, we integrate multiple LLMs to jointly analyze visual, audio, and semantic cues. Their outputs are fused via a scoring mechanism, with final results accepted only when confidence exceeds a high threshold.
Video Scene Segmentation. Video scene segmentation is crucial for dividing video content into coherent semantic units. Traditional methods often rely on low-level visual features, leading to fragmented outputs, especially in dramas with frequent shot transitions. To address this, we propose a three-stage approach: visual shot segmentation, local shot fusion, and global semantic refinement. We first apply a visual segmentation model (e.g., TransNetV2 (Soucek and Lokoc, 2024) and AutoShot Zhu et al. (2023)) to obtain shot-level boundaries. Neighboring shots are then clustered via a trained classifier to reduce fragmentation. Finally, an MLLM refines segmentation globally, merging visually dissimilar but semantically related segments (e.g., close-ups and wide shots of the same scene). This approach balances segmentation granularity and semantic completeness, ensuring transition fluency.
Comprehensive Caption. The Comprehensive Caption module generates holistic semantic summaries of films and dramas. Unlike conventional methods that describe isolated segments, it integrates multimodal inputs from earlier analysis steps to produce coherent, context-aware narratives. Specifically, key information such as characters, dialogues, relationships, and prior episode summaries are fed into an MLLM (e.g., Gemini 1.5 pro Team et al. (2024)) alongside the current video. This ensures accurate, continuous descriptions that capture scene contexts, logical connections, and overall plot development.
Dataset Domain Videos Duration Resolution Reference Edited Videos
@@
DramaAD (Ours) Drama 831 22h 720P-1080P 522
@@
3.2 Human-Inspired Editing Framework
To explore better editing methods, we consult professional editors who engage in commercial editing (e.g., advertisement and promotional video editing) and distill their strategies into our editing framework, which consists of three components: Highlight Detection, Opening & Ending Selection, and Content Pruning. Our automatic editing algorithm is shown in Algorithm 1 as pseudo code.
Highlight Detection. Highlight clips are crucial for producing engaging videos Yao et al. (2016); Xiong et al. (2019); Badamdorj et al. (2022). In our framework, an LLM is introduced to identify highlights. Specifically, we abstract common highlight clips in short dramas and films into a set of rules and use LLM to evaluate whether each scene matches these rules (See Appendix A.3 for more details). Each scene is scored based on the number of matched patterns, while unmatched scenes are labeled as "general scenes" with zero scores.
After scoring, adjacent non-zero-scored scenes are merged into "highlight clips", with the total score reflecting their importance. As illustrated in Figure 2(c), highlight clip 1 consists of adjacent scenes 4, 5, and 6, whose individual scores sum up to 3 + 4 + 2 = 9.
Figure 3: The distribution of the total video duration for each short drama.
Opening & Ending Selection. After highlight clips are located, the next step is to define the scope of the editing. To achieve this, rules can be established to select appropriate start and end scenes for each highlight clip with the top-k scores. For short drama advertisements, the opening aims to attract viewers’ attention, while the ending leaves suspense or concludes climax to encourage clicks.
As highlight clips typically consist of continuous narrative content, we avoid selecting internal scenes within a highlight clip as the opening or ending scenes, in order to maintain the integrity of the storyline. Regarding this, we consider all general scenes preceding the current highlight clip and the first scene of all highlight clips as opening candidates. Similarly, we designate all general scenes following the highlight clip and the final scene of all highlight clips as ending candidates. Take highlight clip 1 in Figure 2(2) as an example, the scenes from 1 to 4 are opening candidates, while scenes 6, 7, 9, and 10 are ending candidates (scene 8 is not the last scene of highlight clip 2).
The LLM evaluates these candidates (details in Appendix A.4) and selects high-quality openings and endings. For each highlight clip, we enumerate all possible combinations of start and end scenes, and then deduplicate on these combinations. This approach enables us to generate diverse editing time ranges.
Content Pruning. Even after boundary selection, some redundant scenes may remain. By removing these redundant scenes, the final video can be made more concise and enhance the viewing experience for the audience. To streamline the final video, we provide the pruning rule (Appendix A.5) for LLM to analyze the relevance between each general scene and the storyline. While ensuring the smooth progression of the video’s storyline, the less captivating scenes are to be pruned. The ideal result is an edited video with a captivating introduction, tightly packed highlights, and a suspenseful conclusion.
4 DramaAD Dataset
4.1 Data Collection
We collect 30 popular Chinese short dramas from the internet, with each drama comprising the first 30% episodes, which is the typical range used for advertisement video editing. Subsequently, we hire some advertisement editing experts to create advertisement videos for these short dramas as references. Additionally, we annotate the reference videos with the corresponding editing timestamps to facilitate future research. In this manner, we construct the proposed dataset, named DramaAD.
@@ 4.2 Statistics and Comparison
DramaAD contains 1353 videos, of which 831 are short drama videos and 522 are edited advertisement videos. These manually edited videos with annotations can serve as golden references for evaluating or fine-tuning the editing model. Although the DramaAD dataset includes only the first 30% of episodes from each short drama series, these videos still span a total duration ranging from 25 minutes to over an hour, as shown in Figure 3. This provides a narrative scale comparable to that of feature-length films.
Compared to other datasets, our drama videos have 1080P resolution (the resolution of reference videos ranges from 720P to 1080P) and are in a 9:16 aspect ratio, which aligns better with the trend of increasing popularity of mobile short-form videos. Additionally, to the best of our knowledge, DramaAD is the first short drama dataset that includes reference videos for advertisement editing.

### v2

@@ Abstract
The rapid growth of online video content, especially on short video platforms, has created a growing demand for efficient video editing techniques that can condense long-form videos into concise and engaging clips. Existing automatic editing methods predominantly rely on textual cues from ASR transcripts and end-to-end segment selection, often neglecting the rich visual context and leading to incoherent outputs. In this paper, we propose a Human-Inspired automatic Video Editing framework (HIVE) that leverages multimodal narrative understanding to address these limitations. Our approach incorporates character extraction, dialogue analysis, and narrative summarization through multimodal large language models, enabling a holistic understanding of the video content. To further enhance coherence, we apply scene-level segmentation and decompose the editing process into three subtasks: highlight detection, opening/ending selection, and pruning of irrelevant content. To facilitate research in this area, we introduce DramaAD, a novel benchmark dataset comprising over 2500 short drama episodes and 500 professionally edited advertisement clips. Experimental results demonstrate that our framework consistently outperforms existing baselines across both general and advertisement-oriented editing tasks, significantly narrowing the quality gap between automatic and human-edited videos.
@@ 1 Introduction
To address these challenges, we propose HIVE, an automatic video editing framework based on multimodal narrative understanding. The HIVE considers both the visual content and the underlying storyline, enabling the generation of diverse and engaging edited videos. Specifically, to achieve accurate video comprehension, we incorporate not only character extraction and dialogue analysis but also leverage MLLMs to generate narrative summaries of video content, thereby compensating for missing visual information. To enhance the coherence and viewing experience of the edited clips, we adopt an improved video scene segmentation Rao et al. (2020), instead of conventional shot detection Fu et al. (2017); Soucek and Lokoc (2024), to divide the video into semantically complete scenes. This ensures that clip boundaries do not interrupt ongoing dialogue or character actions.
@@
To support research in automatic video editing, we present DramaAD, a novel benchmark dataset constructed from real-world short drama videos. These short dramas, fast-paced serialized episodes typically lasting 1 to 3 minutes, have recently surged in popularity on short video platforms. In practical applications, companies are increasingly interested in repurposing such content into advertisement-style clips through automatic editing, aiming to reduce manual effort and production costs. DramaAD addresses this emerging need by offering both academic and industrial value: it comprises over 2500 raw drama videos and more than 500 professionally edited advertisement clips as references. To facilitate objective evaluation, we also propose a suite of quantitative metrics covering multiple aspects of editing quality. Experimental results demonstrate that our proposed framework significantly outperforms existing baselines in both ad-specific and general video editing tasks, while closing the performance gap with human editors.
@@ 2.1 Automatic Video Editing
In recent years, automatic video editing has gained increasing attention Soe (2021); Ronfard (2021); Koorathota et al. (2021); Wang et al. (2019). Early methods primarily relied on feature extraction for video editing. Pardo et al. (2021) trained a ranking model using contrastive learning techniques, which can rank and identify plausible cuts among all possible transitions between a given pair of shots. Podlesnyy (2020) designed a data-driven editing model that treats the video editing task as a sequence decision problem, allowing the model to determine whether each shot should be retained. Hu et al. (2023) utilized a pretrained Vision-Language Model to extract editing-related representations and train an editor model by reinforcement learning. These methods lack an explicit understanding of the video content and suffer from a lack of interpretability. Recently LLM-based editing methods have made progress. Some works Barua et al. (2025); Wang et al. (2024a) utilize LLMs to assist human editors, while other fully automated approaches, such as FunClip, leverage ASR-transcribed speech to achieve video editing.
@@ 2.2 Video Understanding
Recently, significant progress has been made in LLM-based video understanding. Methods such as VideoChat2 (Li et al., 2024) and LLaVA-Video (Lin et al., 2023) integrate video foundation models with LLMs. VideoChat2 uses a learnable Q-former inspired by BLIP2 (Li et al., 2023) for aligning visual and textual features, while LLaVA-Video employs a simple linear projection trained jointly on image-video data. Other methods, such as ChatVideo (Wang et al., 2023a) and VideoAgent (Wang et al., 2024c), incorporate LLMs within agent frameworks. Furthermore, general-purpose multimodal models, such as GPT-4o (Hurst et al., 2024) and Gemini (Team et al., 2023), also demonstrate powerful visual understanding abilities applicable to video tasks. In addition, Some studies focus on understanding films through audio description (Han et al., 2023), narration (Zhang et al., 2024), and storytelling (He et al., 2024), requiring finer video understanding and stronger narrative flow. In the short-video domain, datasets such as SkyScript-100M (Tang et al., 2024) provide structured textual formats capturing scenes, emotions, and cinematography instructions, thus supporting AI-based video generation tasks. Similarly, Vript (Yang et al., 2025) creates detailed mappings between general videos and fine-grained screenplay text, enabling more precise video description and understanding.
In contrast to these works, our research focuses specifically on video semantic analysis and leverages multi-dimensional information to achieve deeper video understanding.
@@ 3 Method
Figure 1: An overview of the proposed framework. The video understanding module (a) extracts video information across multiple dimensions, such as dialogues and characters, etc., and segments the video into semantically complete scenes. Consequently, we obtain a sequence of scenes along with their corresponding narrative descriptions (b). This content is then fed into the editing module (c), which identifies highlights, selects appropriate editing boundaries, and removes unnecessary scenes as human editors do (take highlight clip 1 for example).
@@ 3.1 Multimodal Narrative Understanding
Our framework combines audiovisual models with MLLMs to convert video content into detailed textual representations, simulating human cognition. It features dialogue analysis, character extraction, character-dialogue matching, video scene segmentation and memory module.
Dialogue Analysis. Films and short dramas often feature dense dialogues, with key plot developments driven by character conversations. To capture this, we first apply ASR technology Radford et al. (2023), though its transcripts are prone to errors from acoustic interference, background music, and character name ambiguities. To address these issues, we use optical character recognition (OCR) Bautista and Atienza (2022) to extract subtitles directly from video frames, providing a reliable reference. Leveraging Doubao 1.5 pro  [^3] https://seed.bytedance.com/en/special/doubao_1_5_pro, we align and correct ASR outputs based on OCR subtitles, significantly improving transcription accuracy. Additionally, we perform speaker diarization Horiguchi et al. (2020) to distinguish audio by speaker, enabling fine-grained narrative analysis.
Character Extraction. The goal of character extraction is to identify individuals in videos and build a database of their key information. While annotated data or metadata may sometimes be available, it is often missing and must be generated automatically. To address this, we employ face detection Deng et al. (2019) and clustering: detecting faces in video frames, extracting facial embeddings, and clustering them into distinct identities. Additionally, we enrich character profiles by extracting character’s name or identity from drama dialogues. Integrating these information into the character database can facilitate deeper narrative comprehension.
Character-Dialogue Matching. To enable human-like video understanding, we introduce a specialized module that links identified characters to their dialogues and infers their relationships. After applying speaker diarization, dialogue segments are initially unlabeled. To align the speaker with the spoken content, we input the character’s facial image and dialogue information into Doubao 1.5 pro to reason and infer the speaker’s identity. To ensure accuracy, we employ multiple inference iterations followed by a voting mechanism.
Video Scene Segmentation. Video scene segmentation is crucial for dividing video content into coherent semantic units. Traditional methods often rely on low-level visual features, leading to fragmented outputs, especially in dramas with frequent shot transitions. To achieve this, we first apply AutoShot Zhu et al. (2023) to obtain shot-level boundaries and use Qwen2-VL Wang et al. (2024b) to merge shots with the same site (e.g. restaurant or office). Subsequently, we further utilize Gemini 2.0 flash to merge them belonging to the same event based on contextual information and offer. Meanwhile, a description of the current event is generated. This approach balances segmentation granularity and semantic completeness, ensuring transition fluency.
@@
[…]
@@
8:  $O=\{o_{1},...,o_{m}\}\leftarrow L(R_{o},h^{\prime},V)$ $\triangleright$ Select all suitable opening scenes for $h^{\prime}$, $O\subseteq V$
9:  $E=\{e_{1},...,e_{n}\}\leftarrow L(R_{e},h^{\prime},V)$ $\triangleright$ Select all suitable ending scenes for $h^{\prime}$, $E\subseteq V$
10:  $B\leftarrow B\cup(O\times E)$ $\triangleright$ Enumerate all openings-endings pairs by cartesian product
@@
14:  $P=\{p_{1},...,p_{l}\}\leftarrow L(R_{p},V[o:e])$ $\triangleright$ Get scenes to be pruned within $V[o,e]$, $P\subseteq V$
15:  $V^{\prime}\leftarrow V^{\prime}\cup\text{Cut \& Splice}(V[o:e]\setminus P)$ $\triangleright$ Cut and splice remaining scenes
@@
3.2 Human-Inspired Editing Framework
To explore better editing methods, we consult several professional advertisement editors and distill their strategies into our editing framework, which consists of three components: Highlight Detection, Opening & Ending Selection, and Content Pruning. Algorithm 1 provides a detailed illustration of the workflow of the editing algorithm.
Highlight Detection. Highlight clips are crucial for producing engaging videos Yao et al. (2016); Xiong et al. (2019); Badamdorj et al. (2022). In our framework, DeepSeek-R1 Guo et al. (2025) is introduced to identify highlights. Specifically, we abstract common highlight clips in short dramas and films into a set of rules and use R1 to evaluate whether each scene matches these rules (See Appendix A.3 for more details). Each scene is scored based on the number of matched patterns, while unmatched scenes are labeled as "general scenes" with zero scores. After scoring, adjacent non-zero-scored scenes are merged into "highlight clips", with the total score reflecting their importance.
Opening & Ending Selection. After highlight clips are detected, we select appropriate starting and ending scenes to form the video boundaries. For short drama advertisements, the opening aims to attract viewers’ attention, while the ending leaves suspense or concludes climax to encourage clicks. Candidate starting scenes include preceding general scenes and the first scene of each highlight clip; ending candidates include following general scenes and the last scene of highlight clips. The GPT-4o evaluates these candidates (details in Appendix A.4), and several opening-ending pairs are sampled to ensure variety.
Content Pruning. Even after boundary selection, some redundant scenes may remain. To streamline the final video, the GPT-4o analyzes each general scene’s relevance to the storyline. Less engaging scenes are pruned, ensuring a smooth narrative flow without unnecessary content. The result is an edited video with a captivating introduction, tightly packed highlights, and a suspenseful conclusion.
It is important to emphasize that we only present an editing framework. The framework itself is model-agnostic, allowing any required models to be replaced with more suitable alternatives depending on the practical context. Thus, our framework ensures both flexibility and adaptability across different applications. All prompts used in our framework are listed in Appendix A.
4 DramaAD Dataset
4.1 Data Collection
Firstly, we collect 30 popular Chinese short dramas from the internet. Subsequently, we hire some advertisement editing experts to create advertisement videos for these short dramas as references. Additionally, we annotate the reference videos with the corresponding multimodal narration and editing timestamps to facilitate future research. In this manner, we construct the proposed dataset, named DramaAD.
Dataset Domain Videos Duration Resolution Reference Videos
@@ 4.1 Data Collection
DramaAD (Ours) Drama 2.6K 68h 720P-1080P 522
@@ 4.2 Statistics and Comparison
DramaAD contains 3104 videos, of which 2582 are short drama videos and 522 are edited advertisement videos. Compared to other datasets, our drama videos have 1080P resolution (the resolution of reference videos ranges from 720P to 1080P) and are in a 9:16 aspect ratio, which aligns better with the trend of increasing popularity of mobile short-form videos. The statistical information of the dataset and comparison with other datasets are presented in Figure 2 and Table 1 seperately.
Figure 2: The distribution of the total video duration for each short drama. The average duration is about 40 minutes.
4.3 Evaluation Metrics
To conduct a comprehensive evaluation, we propose multiple quantitative metrics for assessing both the edited videos and the editing methods. Diversity. Diversity is utilized to measure the creativity of the editing methods. Give raw footage, assuming the editing method generates $n$ edited videos $(V_{1},V_{2},...V_{n})$, the diversity is calculated as follows:
Smoothness. Abrupt transitions are a common issue in automated video editing. We use a metric called smoothness to evaluate the overall coherence of the edited video. Smoothness is defined as the average uninterrupted duration, meaning the average number of minutes before an unnatural or jarring transition occurs.
$\mathrm{VEI}=\mathrm{Engagement}\times\mathrm{Smoothness}$ (4)
Besides general metrics, we propose two additional metrics specifically for advertising scenarios.
Hook Rate. It indicates whether viewers will pause to watch the video upon encountering it, reflecting the attractiveness of the video’s opening.
Suspense Rate. It measures whether viewers are motivated to continue watching due to the suspense created at the end of the video.
5 Experiments
5.1 Experiment Settings